# Reparo: Loss-Resilient Generative Codec for Video Conferencing

### v2

@@
Abstract.
@@
Packet loss during video conferencing often leads to poor quality and video freezing. Attempting to retransmit lost packets is often impractical due to the need for real-time playback. Employing Forward Error Correction (FEC) for recovering the lost packets is challenging as it is difficult to determine the appropriate redundancy level. To address these issues, we introduce Reparo — a loss-resilient video conferencing framework based on generative deep learning models. Our approach involves generating missing information when a frame or part of a frame is lost. This generation is conditioned on the data received thus far, taking into account the model’s understanding of how people and objects appear and interact within the visual realm. Experimental results, using publicly available video conferencing datasets, demonstrate that Reparo outperforms state-of-the-art FEC-based video conferencing solutions in terms of both video quality (measured through PSNR, SSIM, and LPIPS) and the occurrence of video freezes.
1. Introduction
@@
Video conferencing applications often suffer from packet loss, resulting in diminished quality and video freezing. This problem is exacerbated by the strong dependence between encoded frames in traditional video codecs. For instance, P-frames depend on previous I-frames, and hence the loss of an I-frame affects many subsequent frames, causing a jarring experience where the video freezes and subsequently exhibits poor quality until the codec can recover from these dependencies.
@@
Existing systems employ two main techniques to combat this problem: retransmission and forward error correction (FEC). Since real-time applications such as video conferencing must recover lost packets within a limited latency to meet the real-time playback requirement, retransmission is only suitable for scenarios with short round trip times. In all other cases, such applications rely on FEC to recover lost packets within acceptable latency. FEC schemes send redundant packets, known as “parity” packets, to recover the lost data using traditional block codes (Reed and Solomon, 1960; MacKay, 2005) or latency-optimized streaming codes (Rudow et al., 2023). However, all FEC approaches face a challenge in choosing how much redundancy to add since losses in the Internet are bursty and unpredictable. Too much redundancy leads to inefficiency and wasted bandwidth, and too little redundancy leads to undecodable frames, causing video freezes and quality degradation in subsequent frames.
@@
This paper presents a novel approach to loss recovery in video conferencing, eschewing the use of redundant packets or retransmission requests. Instead, when a loss occurs, the receiver leverages the power of ”generative” models to reconstruct the missing information. This approach builds upon recent advancements in generative deep learning models (Li et al., 2022, 2023; Liu et al., 2023; Rombach et al., 2022; He et al., 2022), which are capable of reconstructing images (i.e., frames) even when a significant portion of the data is missing (He et al., 2022; Li et al., 2022). Unlike conventional video codecs, which solely rely on received data for frame reconstruction, generative models operate similarly to humans. They utilize a wealth of knowledge about how people and objects appear, move, and interact to generate the missing information. For instance, when presented with one eye of a person, they can generate the missing eye, or with a partial view of an arm, they can reconstruct the entire torso. The key insight here is that such a generative model is ideally suited for a loss-resilient video codec. However, instead of being guided by textual prompts as seen in typical image generators like DALL-E 2, it can be directed by the received pixels in the current frame, as well as information from past frames, to generate the missing content. This guidance ensures that the generated pixels remain compatible with the correctly received data in the current frame and previous frames, eliminating the potential for hallucinating incompatible content, while generating the data in the lost packets.
@@
We introduce Reparo, a loss-resilient generative codec for video conferencing. The design of Reparo involves two steps. In the initial step, it learns to represent the specific video domain of interest, namely video conferencing, using a small dictionary of visual tokens, where each token refers to a patch in a frame. Reparo then operates on these tokens. The transmitter uses a neural network to encode each frame into its respective set of tokens, packetizes them, and transmits the data. Some of these packets may get lost in the network. The Reparo receiver has a neural network that can regenerate the missing tokens from those it receives and its knowledge of how tokens relate to each other in the visual world. Finally, the reconstructed tokens are decoded to produce the original frame. Fig. 1 illustrates the components of Reparo.
@@
Efficient Compression: Reparo efficiently compresses data by capturing prevalent visual features and dependencies among objects and shapes within its dictionary or codebook. This dictionary is pre-learned and known to both the transmitter and receiver, allowing Reparo to transmit only token indices instead of the actual tokens and their underlying dependencies.
Target Bitrate Compliance: Traditional video codecs exhibit variable bit rates due to the significant size discrepancy between I-frames and other frames. This inconsistency makes it challenging for these codecs to meet a precise target bitrate, leading to fluctuations during transmission, transient congestion, and an elevated risk of packet loss or delay. Reparo, in contrast, maintains a constant bitrate as all frames are treated equally, making it easy to adapt to any desired bitrate.
One-Way Communication: Video conferencing frameworks including those that rely on FEC (Rudow et al., 2023; web, [n. d.]) typically rely on the receiver to send an ACK for every decodable frame. The sender waits for the ACK to retransmit, which could cause even longer freezes when the round-trip time is long. In contrast, in Reparo, the receiver does not need to communicate with the transmitter about undecodable frames; it will always use the received tokens to reconstruct the lost ones.
We have conducted an extensive evaluation of Reparo, comparing it with FEC schemes integrated with WebRTC (web, [n. d.]) (ULPFEC and flexFEC), and Tambur (Rudow et al., 2023), a state-of-the-art streaming-code-based FEC approach. The evaluation uses a large corpus of publicly available video clips spanning 5 hours from 84 individuals, which is significantly larger and more diverse than the validation set in prior works on video conferencing (Sivaraman et al., 2022; Rudow et al., 2023; Cheng et al., 2023). The results are as follows:
Reparo consistently improves the visual quality of displayed videos across all loss levels. It achieves 33.4 dB, 32.9 dB, and 31.6 dB for the 10% worst PSNR under low, medium, and high loss levels, respectively, outperforming state-of-the-art VP9+Tambur by 11.5 dB, 16.4 dB, and 14.7 dB. Notably, this loss resilience does not come at the expense of efficient coding; Reparo achieves similar or better PSNR than the baselines in the absence of packet loss.
@@
In summary, traditionally video communication has involved a trade-off between efficiency and resilience. To maximize efficiency, codecs encode frames together (as a delta from a reference frame), whereas to maximize resilience, each frame should be encoded separately. Reparo stands out as the first codec to encode each frame independently, with no reliance on other frames, while maintaining efficiency akin to state-of-the-art video conferencing codecs that encode frames together. We believe that Reparo underscores the potential of interdisciplinary design, marrying advances in computer vision with core principles in coding theory and communication systems.
2. Related Work
@@
Video Codecs. Video applications typically use classical codecs such as VP8, VP9, H.264, H.265, and AV1 (Bankoski et al., 2011; Mukherjee et al., 2015; Schwarz et al., 2007; Sullivan et al., 2012; av1, [n. d.]). These codecs compress video frames using block-based motion prediction, separating them into keyframes (I-frames) that are compressed independently and predicted frames (P-/B-frames) that are compressed based on differences between adjacent frames. While classical codecs are widely supported and efficient in slow modes, in real-time video-conferencing modes, they suffer from the inability to accurately match a desired target bitrate, which leads to packet loss and frame corruption when they exceed available capacity.
@@
To overcome some of these limitations, several neural codecs have been proposed in recent years (Dasari et al., 2022; Yeo et al., 2018; Khani et al., 2021; Wang et al., 2021a; Lu et al., 2019). These codecs use a low-quality video that is then enhanced using a Deep Neural Network (DNN). Reparo differs from such prior neural codecs in two ways. First, these neural codecs still rely on temporal dependencies, wherein an undecodable frame can cause one or more subsequent frames to freeze. In contrast, Reparo has no dependencies between encoded frames and hence the impact of a loss in one frame does not propagate to other frames. Second, none of the neural approaches use generative neural models that synthesize images from a few small pre-computed tokens.
@@
Loss-Resilient Video Codecs. Forward Error Correction (FEC) is a technique used in communication systems to recover lost data packets without retransmission. Instead of retransmission, redundant information sent by the sender is used by the receiver to reconstruct the original data. This is particularly important in real-time communication systems such as VoIP and video conferencing, where retransmission of lost packets can cause unacceptable delays. Traditional FEC codes such as parity codes (Begen, 2010), Reed-Solomon (RS) codes (Reed and Solomon, 1960), and fountain codes (MacKay, 2005) are all block codes that are optimal for random losses, where packets are lost independently. Recently, researchers have proposed using streaming codes for FEC (e.g., Tambur (Rudow et al., 2023)), achieving better loss recovery capabilities than block codes for bursty losses, where several packets over one or more consecutive frames are lost.
@@
Concurrently to our work, GRACE (Cheng et al., 2023) (To appear in NSDI 2024) proposed a neural video codec that can tolerate packet loss. However, there are two key differences with our approach. First, GRACE does not include an explicit loss recovery network. It improves the loss-resilience of an existing neural decoder (DVC (Lu et al., 2019)) by randomly masking information (i.e., using dropout) during its training. Unlike our loss recovery network, GRACE’s decoder cannot leverage multiple received frames to reconstruct the missing information. Second, the DVC code underlying GRACE is based on Delta coding which creates temporal dependencies between encoded frames. Therefore, errors in one decoded frame due to packet loss propagate to subsequent frames even if they incur no loss. On the other hand, Reparo encodes each frame independently, and thus does not have error propagation.
@@
Generative Neural Networks. In recent years, there has been significant progress in the development of generative models, which can create text, audio, images, and videos indistinguishable from those created by humans (Devlin et al., 2018; Liu et al., 2023; Rombach et al., 2022; Li et al., 2022). These models use knowledge of the target domain to generate content under certain conditions. For example, a text generative model can generate a paragraph conditioning on text prompts (Devlin et al., 2018), and an image generative model can produce an image using only a partial view (Li et al., 2022).
@@
To enhance the use of domain knowledge, many recent visual generative models have adopted a two-stage design (van den Oord et al., 2017; Razavi et al., 2019; Chang et al., 2022; Yu et al., 2021; Lee et al., 2022). First, they learn to represent the target domain using a visual token dictionary (i.e., codebook). Each visual token corresponds to a patch in the image and the dictionary serves as a high-level abstraction of the visual world. Generation is then performed in this token space, similar to text generative models. These models have shown impressive performance in image generative tasks, such as text-to-image synthesis (Rombach et al., 2022; Chang et al., 2023) and image editing (Li et al., 2022).
@@
3. Reparo Design
3.1. Overview
@@
3.2. Reparo Components
3.2.1. The Codec
@@
We use a tokenizer called VQGAN (Esser et al., 2021), which consists of an encoder, a decoder, and a codebook (see Fig. 2). The encoder is a convolutional neural network (CNN) that takes patches in an image and maps each one of them to the nearest neighbor vector in the codebook, i.e., the nearest token. The decoder is also a CNN that takes a concatenation of tokens that represent an image and reproduces the original image.
The compression achieved by VQGAN depends on two of its parameters: the number of tokens used for each frame, and the size of the codebook. Since the image is divided into patches, each mapped to a token, the number of tokens dictates the size of each patch within an image. As the number of tokens is increased, the smaller each patch becomes. More tokens allow a more fine-grained reconstruction as it is easier for a token to represent a smaller patch. However, since we transmit token indices from the sender to the receiver, more tokens means more bits for transmitting all of their indices, and reduces the compression factor. Similarly, a larger codebook or dictionary enables a more diverse set of features to choose from for each token, but requires more bits to represent each token index. Thus, both of these parameters lead to different tradeoffs for the achieved bitrate and visual quality. We show this in Fig. 12.
3.2.2. The Packetizer
@@ 3.2.3. The Bitrate Controller
3.2.4. Loss Recovery Module
@@
We use a common neural network architecture called Vision Transformer (ViT) (Dosovitskiy et al., 2021). Transformers have gained widespread popularity in computer vision and natural language processing for predicting missing image patches or words (Devlin et al., 2018; Dosovitskiy et al., 2021; He et al., 2022). The ViT employs an attention module in each layer to aggregate information from all tokens in an image. To predict a missing token, the attention module uses the received tokens and weighs them by their relevance to the missing token. The relevance is computed by performing a softmax over the dot product of each token with every other token. To extend the standard vision transformer structure to video clips, we use a spatio-temporal attention module (Arnab et al., 2021). In each transformer block, we perform attention over the time dimension (across adjacent frames) and then over the space dimension within a frame. This enables our loss recovery module to exploit both spatial information from the same frame and temporal information across consecutive frames. Specifically, to generate a missing token, the module can use the nearby tokens in both space and across frames, as those tokens have a strong correlation with the missing token. Performing attention over time and space sequentially significantly reduces the computational cost: attention over both space and time simultaneously requires of GPU memory, while attention first over time and then over space requires only of GPU memory.
@@
where represents the subset of received tokens in , and is the probability distribution over the codebook for position in the -th frame predicted by the reconstruction network, conditioned on the input received tokens . As is common practice, we only optimize this loss on the missing tokens of the last frame. Optimizing the loss on all tokens reduces reconstruction performance, as previously observed (He et al., 2022). Detailed training schemes are included in §A.
Inference Routine As the deadline for displaying each frame is hit every 33 ms for 30 fps, we aggregate all received packets for the current frame (as identified by the frame and packet indices) and regard all unreceived packets as lost. Once we place the received tokens in their respective positions corresponding to patches in the frame, we can determine the exact locations of the missing tokens. We use all the tokens received from the previous 6 frames to perform spatio-temporal loss recovery.

### v3

@@
Abstract
Packet loss during video conferencing often results in poor quality and video freezing. Retransmitting lost packets is often impractical due to the need for real-time playback, and using Forward Error Correction (FEC) for packet recovery is challenging due to the unpredictable and bursty nature of Internet losses. Excessive redundancy leads to inefficiency and wasted bandwidth, while insufficient redundancy results in undecodable frames, causing video freezes and quality degradation in subsequent frames.
@@ Abstract
We introduce Reparo — a loss-resilient video conferencing framework based on generative deep learning models to address these issues. Our approach generates missing information when a frame or part of a frame is lost. This generation is conditioned on the data received thus far, considering the model’s understanding of how people and objects appear and interact within the visual realm. Experimental results, using publicly available video conferencing datasets, demonstrate that Reparo outperforms state-of-the-art FEC-based video conferencing solutions in terms of both video quality (measured through PSNR, SSIM, and LPIPS) and the occurrence of video freezes.
1 Introduction
@@
Video conferencing applications are a crucial part of modern life. Despite all the advancements, video conferencing applications still suffer from packet loss, resulting in diminished quality and video freezing [7, 60, 44]. This problem is exacerbated by the strong dependence between encoded frames in traditional video codecs [5, 38, 45, 47, 1]. For instance, in traditional codecs, P-frames (Predicted picture) depend on previous I-frames (Intra-coded picture); hence, the loss of an I-frame affects many subsequent frames, causing a jarring experience where the video freezes and subsequently exhibits poor quality until the codec can recover from these dependencies.
@@
Existing systems employ two main techniques to combat this problem: retransmission and forward error correction (FEC). Since real-time applications such as video conferencing must recover lost packets within a limited latency to meet the real-time playback requirement, retransmission is only suitable for scenarios with short round trip times. In all other cases, such applications rely on FEC to recover lost packets within acceptable latency. FEC schemes send redundant packets, known as “parity” packets, to recover the lost data using traditional block codes [42, 37] or latency-optimized streaming codes [44]. However, all FEC approaches face a challenge in choosing how much redundancy to add since losses in the Internet are bursty and unpredictable. Too much redundancy leads to inefficiency and wasted bandwidth, and too little redundancy leads to undecodable frames, causing video freezes and quality degradation in subsequent frames.
@@
This paper presents a novel approach to loss recovery in video conferencing, without using redundant packets or retransmission requests. Instead, when a loss occurs, the receiver leverages the power of "generative" models to reconstruct the missing information. This approach builds upon recent advancements in generative deep learning models [31, 32, 33, 43, 19], which are capable of reconstructing images (i.e., frames) even when a significant portion of the data is missing [19, 31]. Unlike traditional video codecs, which solely rely on received data for frame reconstruction, generative models operate similarly to humans. They utilize a wealth of knowledge about how people and objects appear, move, and interact to generate the missing information. For instance, when presented with one eye of a person, they can generate the missing eye, or with a partial view of an arm, they can reconstruct the entire torso. The key insight here is that such a generative model is ideally suited for a loss-resilient video codec. However, instead of being guided by textual prompts as seen in typical image generators like DALL-E 2, it can be directed by the received pixels in the current frame, as well as information from past frames, to generate the missing content. This guidance ensures that the generated pixels remain compatible with the correctly received data in the current frame and previous frames, eliminating the potential for hallucinating incompatible content, while generating the data in the lost packets.
@@
We introduce Reparo, a loss-resilient generative codec for video conferencing. The design of Reparo involves two steps. In the initial step, it learns to represent the specific video domain of interest, namely video conferencing, using a small codebook of visual tokens, where each token refers to a patch in a frame. Reparo then operates on these tokens. The transmitter uses a neural network to encode each frame into its respective set of tokens, packetizes them, and transmits the data. Some of these packets may get lost in the network. The Reparo receiver has a neural network that can regenerate the missing tokens from those it receives and its knowledge of how tokens relate to each other in the visual world. Finally, the reconstructed tokens are decoded to produce the original frame. Fig. 1 illustrates the components of Reparo.
@@
1.
Efficient Compression: Reparo efficiently compresses data by capturing prevalent visual features and dependencies among objects and shapes within its codebook. This codebook is pre-learned and known to both the transmitter and receiver, allowing Reparo to transmit only token indices instead of the actual tokens and their underlying dependencies.
2.
Target Bitrate Compliance: Traditional video codecs exhibit variable bitrates partially due to the significant size discrepancy between I-frames and other frames. This inconsistency makes it challenging for these codecs to meet a precise target bitrate, leading to fluctuations during transmission, transient congestion, and an elevated risk of packet loss or delay. Reparo, in contrast, maintains a constant bitrate as all frames are treated equally, making it easy to adapt to any desired bitrate.
3.
One-Way Communication: Video conferencing frameworks, including those that rely on FEC [44, 3], typically rely on the receiver to send an ACK for every decodable frame. The sender waits for the ACK to retransmit, which could cause even longer freezes when the round-trip time is long. In contrast, in Reparo, the receiver does not need to communicate with the transmitter about undecodable frames; it will always use the received tokens to reconstruct the lost ones.
We have conducted an extensive evaluation of Reparo, comparing it with FEC schemes integrated with WebRTC [3] (ULPFEC and flexFEC), and Tambur [44], a state-of-the-art streaming-code-based FEC approach. The evaluation uses a large corpus of publicly available video clips spanning 5 hours from 84 individuals, which is significantly larger and more diverse than the validation set in prior works on video conferencing [46, 44, 11]. The results are as follows:
1.
Reparo consistently improves the visual quality of displayed videos across all loss levels. It achieves 33.4 dB, 32.9 dB, and 31.6 dB for the 10% worst PSNR (Peak Signal to Noise Ratio) values under low, medium, and high loss levels, respectively, outperforming state-of-the-art integration of VP9 (a classical video codec)+Tambur by 11.5 dB, 16.4 dB, and 14.7 dB. Notably, this loss resilience does not come at the expense of efficient coding; Reparo achieves similar or better PSNR compared to baselines in the absence of packet loss.
2.
@@
In summary, video communication has involved a trade-off between efficiency and resilience traditionally. To maximize efficiency, codecs encode frames together (as a delta from a reference frame), whereas to maximize resilience, each frame should be encoded separately. Reparo stands out as the first codec to encode each frame independently, with no reliance on other frames, while maintaining efficiency akin to state-of-the-art video conferencing codecs that encode frames together. We believe that Reparo underscores the potential of interdisciplinary design, marrying advances in computer vision with core principles in coding theory and communication systems.
2 Related Work
@@
Video Codecs. Video applications typically use classical codecs such as VP8, VP9, H.264, H.265, and AV1 [5, 38, 45, 47, 1]. These codecs compress video frames using block-based motion prediction, separating them into keyframes (I-frames) that are compressed independently and predicted frames (P-/B-frames) that are compressed based on differences between adjacent frames. While classical codecs are widely supported and efficient in slow modes, in real-time video conferencing modes, they are unable to accurately match a desired target bitrate, which leads to packet loss and frame corruption when they exceed available capacity [24].
@@
To overcome some of these limitations, several neural codecs have been proposed in recent years [12, 56, 28, 53, 36]. These codecs use a low-quality video that is then enhanced using a Deep Neural Network (DNN). Reparo differs from such prior neural codecs in two ways. First, these neural codecs still rely on temporal dependencies, wherein an undecodable frame can cause one or more subsequent frames to freeze. In contrast, Reparo has no dependencies between encoded frames and hence the impact of a loss in one frame does not propagate to other frames. Second, none of the neural approaches use generative neural models that synthesize images from a few small pre-computed tokens.
@@
Loss-Resilient Video Codecs. Forward Error Correction (FEC) is a technique used in communication systems to recover lost data packets without retransmission. Instead of retransmission, redundant information sent by the sender is used by the receiver to reconstruct the original data. This is particularly important in real-time communication systems such as video conferencing, where retransmission of lost packets can cause unacceptable delays. Traditional FEC codes such as parity codes [6], Reed-Solomon (RS) codes [42], and fountain codes [37] are all block codes that are optimal for random losses, where packets are lost independently. Recently, researchers have proposed using streaming codes for FEC (e.g., Tambur [44]), achieving better loss recovery capabilities than block codes for bursty losses, where several packets over one or more consecutive frames are lost.
@@
GRACE [11] proposed a neural video codec that can tolerate packet loss. However, there are two key differences with our approach. First, GRACE does not include an explicit loss recovery network. It improves the loss-resilience of an existing neural decoder (DVC [36]) by randomly masking information (i.e., using dropout) during its training. Unlike our loss recovery network, GRACE’s decoder cannot leverage multiple received frames to reconstruct the missing information. Second, the DVC code underlying GRACE is based on Delta coding which creates temporal dependencies between encoded frames. Therefore, errors in one decoded frame due to packet loss propagate to subsequent frames even if they incur no loss. On the other hand, Reparo encodes each frame independently, and thus does not have error propagation.
@@
Generative Neural Networks. In recent years, there has been significant progress in the development of generative models, which can create text, audio, images, and videos indistinguishable from those created by humans [13, 33, 43, 31]. These models use knowledge of the target domain to generate content under certain conditions. For example, a text generative model can generate a paragraph conditioning on text prompts [13], and an image generative model can produce an image using only a partial view [31].
@@
To enhance the use of domain knowledge, many recent visual generative models have adopted a two-stage design [51, 41, 10, 57, 30]. First, they learn to represent the target domain using a visual token codebook. Each visual token corresponds to a patch in the image and the codebook serves as a high-level abstraction of the visual world. Generation is then performed in this token space, similar to text generative models. These models have shown impressive performance in image generative tasks, such as text-to-image synthesis [43, 9] and image editing [31].
@@
3 Reparo Design
3.1 Overview
@@
3.2 Reparo Components
3.2.1 The Neural Codec: Encoder and Decoder
@@
We use a tokenizer called VQGAN [16], which consists of an encoder, a decoder, and a codebook (see Fig. 2). The encoder is a convolutional neural network (CNN) that takes patches in an image and maps each one of them to the nearest neighbor vector in the codebook, i.e., the nearest token. The decoder is also a CNN that takes a concatenation of tokens that represent an image and reproduces the original image.
The compression achieved by VQGAN depends on two of its parameters: the number of tokens used for each frame, and the size of the codebook. Since the image is divided into patches, each mapped to a token, the number of tokens dictates the size of each patch within an image. As the number of tokens is increased, the smaller each patch becomes. More tokens allow a more fine-grained reconstruction as it is easier for a token to represent a smaller patch. However, since we transmit token indices from the sender to the receiver, more tokens means more bits for transmitting all of their indices, and reduces the compression factor. Similarly, a larger codebook enables a more diverse set of features to choose from for each token, but requires more bits to represent each token index. Thus, both of these parameters lead to different tradeoffs for the achieved bitrate and visual quality. We show this in Fig. 12.
3.2.2 The Packetizer
@@ 3.2.3 The Bitrate Controller
3.2.4 Loss Recovery Module
@@
We use a common neural network architecture called Vision Transformer (ViT) [14]. Transformers have gained widespread popularity in computer vision and natural language processing for predicting missing image patches or words [13, 14, 19]. The ViT employs an attention module in each layer to aggregate information from all tokens in an image. To predict a missing token, the attention module uses the received tokens and weighs them by their relevance to the missing token. The relevance is computed by performing a softmax over the dot product of each token with every other token. To extend the standard vision transformer structure to video clips, we use a spatio-temporal attention module [4]. In each transformer block, we perform attention over the time dimension (across adjacent frames) and then over the space dimension within a frame. This enables our loss recovery module to exploit both spatial information from the same frame and temporal information across consecutive frames. Specifically, to generate a missing token, the module can use the nearby tokens in both space and across frames, as those tokens have a strong correlation with the missing token. Performing attention over time and space sequentially significantly reduces the computational cost: attention over both space and time simultaneously requires of GPU memory, while attention first over time and then over space requires only of GPU memory.
@@
where represents the subset of received tokens in , and is the probability distribution over the codebook for position in the -th frame predicted by the reconstruction network, conditioned on the input received tokens . As is common practice, we only optimize this loss on the missing tokens of the last frame. Optimizing the loss on all tokens reduces reconstruction performance, as previously observed [19]. Detailed training schemes are included in §A.
Inference Routine As the deadline for displaying each frame is hit every 33 ms for 30 fps, we aggregate all received packets for the current frame (as identified by the frame and packet indices) and regard all unreceived packets as lost. Once we place the received tokens in their respective positions corresponding to patches in the frame, we can determine the exact locations of the missing tokens. We use all the tokens received from the previous frames to perform spatio-temporal loss recovery.